v> Dr Mukonda has two cars, which he hires out day by day. The number of

demands follows poisson with mean 1.5. Calculate the probability that on

a day: (a) Neither car is used (b) Some demands is refused.


1
Expert's answer
2022-02-07T16:52:58-0500

(a) "P(X=0)=\frac{1.5^0e^{-1.5}}{0!}=0.2231."


(b) "P(X>2)=1-P(X=0)-P(X=1)-P(X=2)="

"=1-e^{-1.5}(\frac{1.5^0}{0!}+\frac{1.5^1}{1!}+\frac{1.5^2}{2!})=0.1912."